# Finding Most Shattering Minimum Vertex Cuts of Polylogarithmic Size in Near-Linear Time

### v1

@@ 5.2 Algorithm and Runtime
At a high level, Algorithm˜1 works as follows. Let $(x,\nu,\Pi)$ be a tuple given as input to the algorithm. For each path $\pi\in\Pi$, we can traverse the path from $x$ to the far-most endpoint of $\pi$. At a given vertex $u\in\pi$, we can explore the bridges of $\Pi$ attached to $u$ using a breadth-first search (BFS). While doing so, we maintain a list of bridges of $\Pi$. We also maintain a list of candidates by checking whether the attachment set of a bridge forms a $k$-tuple. Among the list of candidates, false candidates are then pruned correctly and efficiently via Lemma˜5.3 and Lemma˜4.9. So far, we have not deviated from $\texttt{Shredders}(\cdot,\cdot)$ (see Figure˜3).
@@
However, our modification imposes a restriction on the number of edges we are allowed to explore via BFS. We can explore at most $\nu$ edges along each path. Suppose we are processing a vertex $u$ along path $\pi$. If the number of edges explored for this path exceeds $\nu$ while exploring bridges of $\Pi$ attached to $u$, we terminate early and mark $u$ as an unverified vertex. Intuitively, the unverified vertex $u$ serves as a boundary of exploration. It signifies that bridges attached to $u$ were not fully explored, so we flag $u$ and treat it with extra caution during a later step. At the end of processing all the paths in $\Pi$, we will have obtained a list of bridges of $\Pi$, a list of candidate $k$-shredders of $\Pi$, and an unverified set consisting of all the unverified vertices. The final step is to black box Lemma˜4.9 to prune false candidates among the list of candidate $k$-shredders. Consider Algorithm˜1.
@@
1: $x$ - a sample vertex $\nu$ - volume parameter $\Pi$ - a set of $k$ openly disjoint paths starting from $x$
2: $\mathcal{L}$ - a list of $k$-shredders captured by $(x,\nu,\Pi)$ $U$ - an unverified set
3: $\mathcal{L}\leftarrow\varnothing$
4: $U\leftarrow\varnothing$
5: for each path $\pi\in\Pi$ do
6:   reset explored edges and vertices to null
7:   for each vertex $u\in\pi$ in increasing distance from $x$ do
8:    if $u$ is the far-most endpoint of $\pi$ then
9:      $U\leftarrow U\cup\{u\}$
10:      break
11:    explore all unexplored bridges of $\Pi$ attached to $u$ with BFS: 12:  terminate early the moment more than $\nu$ edges are explored
13:    if BFS terminated early then
14:      $U\leftarrow U\cup\{u\}$
15:      break
16:    else
17:      for each bridge $\Gamma$ explored by BFS do
18:       if the attachment set $A$ of $\Gamma$ forms a $k$-tuple of $\Pi$ then
19:         $\mathcal{L}\leftarrow\mathcal{L}\cup\{A\}$
20: if $x\in U$ then
21:   return $(\varnothing,\varnothing)$
22:
23: prune false $k$-shredders in $\mathcal{L}\cup\{U\}$ using Lemma˜4.9
24: $F\leftarrow$ the set of far-most endpoints of all paths in $\Pi$
25: if $U$ was not pruned and $U\cap F=\varnothing$ then
26:   return $(\mathcal{L},U)$
27: else
28:   return $(\mathcal{L},\varnothing)$
@@
A few lines of Algorithm˜1 were added to handle some special cases. Specifically, lines 8, 20, and 25. The justification of these lines will become clear in the following pages, but we will provide basic reasoning here.
@@
In our case, however, each path starts from $x$ but can end at an arbitrary vertex. There is no common endpoint $y$. The fix for this is the inclusion of line 8. If we reach an endpoint of a path, we should immediately mark the vertex as unverified and consider the endpoints later. Intuitively, because $(x,\nu,\Pi)$ captures $S$, we do not need to explore the far-most endpoints of $\Pi$ because they live in $C$. After all, our goal is to list $S$, which does not contain any of the path endpoints.
Line 20 covers a trivial case. What may occur is that the bridges of $\Pi$ attached to $x$ have total volume greater than $\nu$. In this case, we will terminate the BFS due to line 12 when processing $x$ and never find a candidate $k$-shredder. Right before line 20 we will have $U=\{x\}$. In this case, we can terminate immediately and return empty sets because no candidate $k$-shredders were found. We should think of this case as a “bad input” tuple scenario where $(x,\nu,\Pi)$ did not capture anything.
Line 25 covers the case where the unverified set contains an endpoint of some path in $\Pi$. We need not report $U$ in this case because any $k$-shredder $S$ that $(x,\nu,\Pi)$ captures must use vertices strictly preceding each far-most endpoint of all paths. This may seem like an arbitrary restriction, but it will be useful to include line 25 in the proof of Lemma˜5.11.
@@
Now we analyze the rest of the algorithm. As argued above, the total number of edges we explore over all paths is at most $k\nu$. Since each candidate $k$-shredder corresponds with a distinct bridge of $\Pi$ whose attachments form a $k$-tuple, there can be at most $\frac{k\nu}{k}=\mathcal{O}(\nu)$ candidate $k$-shredders in $\mathcal{L}$. Let $\mathcal{B}$ denote the set of all bridges explored. Any bridge of $\Pi$ at the very least requires one edge, so the volume of all bridges put together is at most $\mathcal{O}(k\nu)$. Since the sum of lengths over all paths is at most $k^{2}\nu$, Lemma˜4.9 implies we can implement line 23 in $\mathcal{O}(k|\mathcal{L}|\log|\mathcal{L}|+k^{2}\nu+\mathrm{vol}(\mathcal{B}))=\mathcal{O}(k\nu\log\nu+k^{2}\nu+k\nu)=\mathcal{O}(k^{2}\nu\log\nu)$ time. ∎
@@ 5.3 Correctness
Now we prove the correctness portion of Lemma˜2.5. These lemmas effectively streamline the argument of [CT99] and show that the structural properties of $\texttt{Shredders}(\cdot,\cdot)$ extend to our localized setting. We will prove two core lemmas.
@@
Let $S$ be a $k$-shredder captured by $(x,\nu,\Pi)$. Then, Algorithm˜1 will either identify $S$ as a candidate $k$-shredder or identify $S$ as the unverified set. Specifically, either $S\in\mathcal{L}$ or $S=U$ by the end of the algorithm.
@@
All sets in $\mathcal{L}$ at the end of Algorithm˜1 are $k$-shredders.
It is straightforward to check that these two lemmas directly imply the algorithm correctness portion of Lemma˜2.5.
@@ 5.3.1 Identification of Captured Shredders
First we prove Lemma˜5.5. Let $S$ be a $k$-shredder captured by $(x,\nu,\Pi)$. We will prove that the algorithm will walk along each path and reach the vertices of $S$. Then, we prove that $S$ will either be marked as a candidate $k$-shredder or the unverified set.
@@
Fix a path $\pi\in\Pi$. Let $P=\{v\in\pi\mid\delta_{\pi}(v)<\delta_{\pi}(\pi(S))\}$. It suffices to prove that the algorithm will not terminate early when exploring bridges attached to $P$. Let $Q_{x}$ denote the component of $\mathcal{R}$ containing $x$. Every vertex in $P$ must be in $Q_{x}$ because they precede $\pi(S)$ along $\pi$. Because all vertices of $S$ are in $\Pi$, we cannot reach a component of $\mathcal{R}\setminus\{Q_{x}\}$ when exploring bridges of $\Pi$ attached to $P$. It follows that all edges explored for the vertices in $P$ must contain an endpoint in $Q_{x}$. Because $Q_{x}\in\mathcal{R}$ we have $\mathrm{vol}(Q_{x})\leq\mathrm{vol}(\mathcal{R})\leq\nu$. Line 12 ensures the algorithm will not terminate early. This means that for all paths $\pi\in\Pi$, the unverified vertex $u\in\Pi$ satisfies $\delta_{\pi}(\pi(S))\leq\delta_{\pi}(u)$. Hence, $S\preceq U$. ∎
Proof of Lemma˜5.5.
Lemma˜5.2 implies that there exists a component $Q$ of $G\setminus S$ that is also a component of $G\setminus\Pi$. This means that $Q$ is a bridge of $\Pi$ whose attachment set is $S$. From Lemma˜5.7 we know that there are two cases. Either $\delta_{\pi}(\pi(S))=\delta_{\pi}(\pi(U))$ for all paths $\pi\in\Pi$, or $\delta_{\pi}(\pi(S))<\delta_{\pi}(\pi(U))$ for some path $\pi\in\Pi$. In the former case, we have that $S=U$. In the latter case, line 11 and lines 17-19 imply we must have fully explored all bridges of $\Pi$ attached to a vertex of $S$. Because $Q$ is a bridge of $\Pi$ attached to all vertices of $S$, we will put $N(Q)=S$ as a set in $\mathcal{L}$. Lemma˜5.3 implies that no bridge of $\Pi$ straddles $S$, which means $S$ will not be pruned from $\mathcal{L}$ on line 23. ∎
@@ 5.3.2 Algorithm Validity
So far we have proved that all $k$-shredders captured by $(x,\nu,\Pi)$ will be returned by Algorithm˜1. However, this is not enough to show correctness; we also need to show that every set returned by the algorithm is a $k$-shredder. To do this, we will first prove that if a set $S\in\mathcal{L}$ is not straddled by a bridge of $\Pi$, then it is a $k$-shredder. Then, we will prove that straddled sets in $\mathcal{L}$ are pruned on line 23. The following fact is trivial but useful for the next two lemmas.
@@
Let $S$ be a set remaining in $\mathcal{L}$ after line 23. Then, there exists a path $\pi\in\Pi$ such that $\delta_{\pi}(\pi(S))<\delta_{\pi}(u)$, where $u$ is the unverified vertex along $\pi$.
@@
Since $S$ was put in $\mathcal{L}$, there must exist a vertex $s\in S$ such that the algorithm explored all bridges of $\Pi$ attached to $s$. This is guaranteed by lines 14 and 16. This implies that $s$ precedes the unverified vertex along its path. ∎
Next we prove that candidates that are not straddled by any bridges of $\Pi$ are $k$-shredders. This can be viewed as a local version of one direction of Lemma˜4.8.
@@
Let $S$ be a set put in $\mathcal{L}$ at some point on line 19. If no bridge of $\Pi$ straddles $S$, then $S$ is a $k$-shredder.
@@
We prove $S$ is a $k$-shredder by showing $G\setminus S$ has at least three components. Firstly, we have $S=N(Q)$ for some component $Q$ of $G\setminus\Pi$. Fix a path $\pi\in\Pi$ such that the vertex $\pi(S)$ is not the far-most endpoint of $\pi$. Such a path must exist by ˜5.8. Let $z$ denote the far-most endpoint of $\pi$. It suffices to show that $x$ is not connected to $z$ in $G\setminus S$. Then, because $x\notin Q$ and $z\notin Q$, there are at least three components of $G\setminus S$: the component containing $x$, the component containing $z$, and $Q$.
@@
Consider the open subpath $P(u{\rightsquigarrow}v)$ (see Figure˜4). We claim that every vertex in $P(u{\rightsquigarrow}v)$ is not in $\Pi$. To see this, suppose that there exists a vertex $w\in P(u{\rightsquigarrow}v)$ such that $w\in\pi_{3}$ for some path $\pi_{3}\in\Pi$. Then, $\delta_{\pi_{3}}(w)=\delta_{\pi_{3}}(\pi_{3}(S))$ because otherwise, it would contradict the definition of $u$ or $v$. This implies $w=\pi_{3}(S)$ because the vertex distance $\delta_{\pi_{3}}(\pi_{3}(S))$ from $x$ along $\pi_{3}$ is unique. However, this contradicts the definition of $P$ because $P$ is a path in $G\setminus S$: it cannot contain $\pi_{3}(S)$.
@@
Let $S$ be a candidate $k$-shredder detected by Algorithm˜1 on input $(x,\nu,\Pi)$. If there exists a bridge of $\Pi$ that straddles $S$, then $S$ will be pruned from the list of candidates. Similarly, if there exists a bridge that straddles the unverified set $U$, then $U$ will be pruned.
@@
Let $S$ be the candidate straddled by a bridge $\Gamma$ of $\Pi$. For a visual example of this proof, see Figure˜5. Let $A$ denote the attachment set of $\Gamma$. If any vertex in $A$ has been fully processed without terminating early, then $\Gamma$ will be detected on line 11. In this case, $S$ will be pruned by line 23.
Hence, we can assume that no vertex of $A$ has been fully explored, i.e. $U\preceq A$. We will show that $U$ straddles $S$, in which case $S$ will be pruned by line 23. First, by ˜5.8, there must exist a path $\pi_{1}\in\Pi$ that satisfies $\delta_{\pi_{1}}(\pi_{1}(S))<\delta_{\pi_{1}}(\pi_{1}(U))$. Because $\Gamma$ straddles $S$, there must exist a path $\pi_{2}\in\Pi$ and a vertex $v\in A\cap\pi_{2}$ such that $\delta_{\pi_{2}}(v)<\delta_{\pi_{2}}(\pi_{2}(S))$. Furthermore, because $U\preceq A$, this implies that $\delta_{\pi_{2}}(\pi_{2}(U))\leq\delta_{\pi_{2}}(v)<\delta_{\pi_{2}}(\pi_{2}(S))$. Now, we have $\delta_{\pi_{1}}(\pi_{1}(S))<\delta_{\pi_{1}}(\pi_{1}(U))$ and $\delta_{\pi_{2}}(\pi_{2}(U))<\delta_{\pi_{2}}(\pi_{2}(S))$, which implies $U$ straddles $S$. Line 23 implies that $S$ will be pruned from the list of candidates.
Figure 5: The bridge $Q$ has attachment set $A=S=\{s_{1},s_{2},s_{3}\}$. $S$ is straddled by $\Gamma$ with attachments $\{u,v\}$. If we explore all bridges of $\Pi$ attached to $u$, then $\Gamma$ will be naturally detected (top). Otherwise, we must have terminated early at some vertex $u_{1}$ before reaching $u$. ˜5.8 implies all bridges of $\Pi$ attached to some vertex $s_{2}$ were explored without terminating early. So $s_{2}$ must precede some unverified vertex $u_{2}$. This implies that $S$ is straddled by the unverified vertices (bottom).
Next, we prove the second statement. Suppose the unverified set $U$ is straddled by a bridge. First, we may assume that $|U|=k$ as otherwise $U=\{x\}$, which is caught by line 20. Suppose that a bridge $\Gamma$ of $\Pi$ straddles $U$ with attachment set $A$. If any vertex in $A$ has been fully explored by the algorithm, then $\Gamma$ will be detected by the BFS and $U$ will be pruned by line 23. Otherwise, we have $U\preceq A$, which contradicts our assumption that $\Gamma$ straddles $U$. ∎
Proof of Lemma˜5.6.
Lemma˜5.9 shows that the sets in $\mathcal{L}$ which are not straddled by a bridge of $\Pi$ are $k$-shredders. Lemma˜5.10 shows that sets in $\mathcal{L}$ (along with the unverified set) which are straddled are removed from $\mathcal{L}$. This implies that at the end of Algorithm˜1, every set in $\mathcal{L}$ is a $k$-shredder. ∎
Finally, we can combine Lemmas˜5.5, 5.6 and 5.4 to prove Lemma˜2.5.
Proof of Lemma˜2.5.
Let $S$ be a $k$-shredder captured by $(x,\nu,\Pi)$. Lemma˜5.5 implies that $S\in\mathcal{L}$ or $S=U$ by the end of the algorithm. Furthermore, every set in $\mathcal{L}$ by the end of the algorithm is a $k$-shredder by Lemma˜5.6. Finally, Lemma˜5.4 proves the time complexity of the algorithm. ∎

### v2

@@ 5.2 Algorithm and Runtime
At a high level, Algorithm 1 works as follows. Let $(x,\nu,\Pi)$ be a tuple given as input to the algorithm. For each path $\pi\in\Pi$, we can traverse the path from $x$ to the far-most endpoint of $\pi$. At a given vertex $u\in\pi$, we can explore the bridges of $\Pi$ attached to $u$ using a breadth-first search (BFS). While doing so, we maintain a list of bridges of $\Pi$. We also maintain a list of candidates by checking whether the attachment set of a bridge forms a $k$-tuple. Among the list of candidates, false candidates are then pruned correctly and efficiently via Lemma 5.3 and Lemma 4.9. So far, we have not deviated from $\texttt{Shredders}(\cdot,\cdot)$ (see Figure 3).
@@
However, our modification imposes a restriction on the number of edges we are allowed to explore via BFS. We can explore at most $\nu$ edges along each path. Suppose we are processing a vertex $u$ along path $\pi$. If the number of edges explored for this path exceeds $\nu$ while exploring bridges of $\Pi$ attached to $u$, we terminate early and mark $u$ as an unverified vertex. Intuitively, the unverified vertex $u$ serves as a boundary of exploration. It signifies that bridges attached to $u$ were not fully explored, so we flag $u$ and treat it with extra caution during a later step. At the end of processing all the paths in $\Pi$, we will have obtained a list of bridges of $\Pi$, a list of candidate $k$-shredders of $\Pi$, and an unverified set consisting of all the unverified vertices. The final step is to black box Lemma 4.9 to prune false candidates among the list of candidate $k$-shredders. Consider Algorithm 1.
@@
$x$- a sample vertex
1:$\nu$ volume parameter
2:$\Pi$ a set of $k$ openly disjoint paths starting from $x$$\mathcal{L}$ a list of $k$-shredders captured by $(x,\nu,\Pi)$
3:$U$ an unverified set
$\mathcal{L}\leftarrow\varnothing$
5:$U\leftarrow\varnothing$Foreach path $\pi\in\Pi$
[…]
7:$U\leftarrow U\cup\{u\}$
[…]
9:explore all unexplored bridges of $\Pi$ attached to $u$ with BFS:
[…]
11:$U\leftarrow U\cup\{u\}$
[…]
15:
16:prune false $k$-shredders in $\mathcal{L}\cup\{U\}$ using Lemma 4.9
[…]
18:return $(\mathcal{L},U)$Else
19:return $(\mathcal{L},\varnothing)$EndIf
[…]
@@
A few lines of Algorithm 1 were added to handle some special cases. Specifically, lines 6, 13, and 17. The justification of these lines will become clear in the following pages, but we will provide basic reasoning here.
@@
In our case, however, each path starts from $x$ but can end at an arbitrary vertex. There is no common endpoint $y$. The fix for this is the inclusion of line 6. If we reach an endpoint of a path, we should immediately mark the vertex as unverified and consider the endpoints later. Intuitively, because $(x,\nu,\Pi)$ captures $S$, we do not need to explore the far-most endpoints of $\Pi$ because they live in $C$. After all, our goal is to list $S$, which does not contain any of the path endpoints.
Line 13 covers a trivial case. What may occur is that the bridges of $\Pi$ attached to $x$ have total volume greater than $\nu$. In this case, we will terminate the BFS due to line 10 when processing $x$ and never find a candidate $k$-shredder. Right before line 13 we will have $U=\{x\}$. In this case, we can terminate immediately and return empty sets because no candidate $k$-shredders were found. We should think of this case as a “bad input” tuple scenario where $(x,\nu,\Pi)$ did not capture anything.
Line 17 covers the case where the unverified set contains an endpoint of some path in $\Pi$. We need not report $U$ in this case because any $k$-shredder $S$ that $(x,\nu,\Pi)$ captures must use vertices strictly preceding each far-most endpoint of all paths. This may seem like an arbitrary restriction, but it will be useful to include line 17 in the proof of Lemma 5.11.
@@
Now we analyze the rest of the algorithm. As argued above, the total number of edges we explore over all paths is at most $k\nu$. Since each candidate $k$-shredder corresponds with a distinct bridge of $\Pi$ whose attachments form a $k$-tuple, there can be at most $\frac{k\nu}{k}=\mathcal{O}(\nu)$ candidate $k$-shredders in $\mathcal{L}$. Let $\mathcal{B}$ denote the set of all bridges explored. Any bridge of $\Pi$ at the very least requires one edge, so the volume of all bridges put together is at most $\mathcal{O}(k\nu)$. Since the sum of lengths over all paths is at most $k^{2}\nu$, Lemma 4.9 implies we can implement line 16 in $\mathcal{O}(k|\mathcal{L}|\log|\mathcal{L}|+k^{2}\nu+\mathrm{vol}(\mathcal{B}))=\mathcal{O}(k\nu\log\nu+k^{2}\nu+k\nu)=\mathcal{O}(k^{2}\nu\log\nu)$ time. ∎
@@ 5.3 Correctness
Now we prove the correctness portion of Lemma 2.5. These lemmas effectively streamline the argument of [CT99] and show that the structural properties of $\texttt{Shredders}(\cdot,\cdot)$ extend to our localized setting. We will prove two core lemmas.
@@
Let $S$ be a $k$-shredder captured by $(x,\nu,\Pi)$. Then, Algorithm 1 will either identify $S$ as a candidate $k$-shredder or identify $S$ as the unverified set. Specifically, either $S\in\mathcal{L}$ or $S=U$ by the end of the algorithm.
@@
All sets in $\mathcal{L}$ at the end of Algorithm 1 are $k$-shredders.
It is straightforward to check that these two lemmas directly imply the algorithm correctness portion of Lemma 2.5.
@@ 5.3.1 Identification of Captured Shredders
First we prove Lemma 5.5. Let $S$ be a $k$-shredder captured by $(x,\nu,\Pi)$. We will prove that the algorithm will walk along each path and reach the vertices of $S$. Then, we prove that $S$ will either be marked as a candidate $k$-shredder or the unverified set.
@@
Fix a path $\pi\in\Pi$. Let $P=\{v\in\pi\mid\delta_{\pi}(v)<\delta_{\pi}(\pi(S))\}$. It suffices to prove that the algorithm will not terminate early when exploring bridges attached to $P$. Let $Q_{x}$ denote the component of $\mathcal{R}$ containing $x$. Every vertex in $P$ must be in $Q_{x}$ because they precede $\pi(S)$ along $\pi$. Because all vertices of $S$ are in $\Pi$, we cannot reach a component of $\mathcal{R}\setminus\{Q_{x}\}$ when exploring bridges of $\Pi$ attached to $P$. It follows that all edges explored for the vertices in $P$ must contain an endpoint in $Q_{x}$. Because $Q_{x}\in\mathcal{R}$ we have $\mathrm{vol}(Q_{x})\leq\mathrm{vol}(\mathcal{R})\leq\nu$. Line 10 ensures the algorithm will not terminate early. This means that for all paths $\pi\in\Pi$, the unverified vertex $u\in\Pi$ satisfies $\delta_{\pi}(\pi(S))\leq\delta_{\pi}(u)$. Hence, $S\preceq U$. ∎
Proof of Lemma 5.5.
Lemma 5.2 implies that there exists a component $Q$ of $G\setminus S$ that is also a component of $G\setminus\Pi$. This means that $Q$ is a bridge of $\Pi$ whose attachment set is $S$. From Lemma 5.7 we know that there are two cases. Either $\delta_{\pi}(\pi(S))=\delta_{\pi}(\pi(U))$ for all paths $\pi\in\Pi$, or $\delta_{\pi}(\pi(S))<\delta_{\pi}(\pi(U))$ for some path $\pi\in\Pi$. In the former case, we have that $S=U$. In the latter case, line 9 and lines 12-13 imply we must have fully explored all bridges of $\Pi$ attached to a vertex of $S$. Because $Q$ is a bridge of $\Pi$ attached to all vertices of $S$, we will put $N(Q)=S$ as a set in $\mathcal{L}$. Lemma 5.3 implies that no bridge of $\Pi$ straddles $S$, which means $S$ will not be pruned from $\mathcal{L}$ on line 16. ∎
@@ 5.3.2 Algorithm Validity
So far we have proved that all $k$-shredders captured by $(x,\nu,\Pi)$ will be returned by Algorithm 1. However, this is not enough to show correctness; we also need to show that every set returned by the algorithm is a $k$-shredder. To do this, we will first prove that if a set $S\in\mathcal{L}$ is not straddled by a bridge of $\Pi$, then it is a $k$-shredder. Then, we will prove that straddled sets in $\mathcal{L}$ are pruned on line 16. The following fact is trivial but useful for the next two lemmas.
@@
Let $S$ be a set remaining in $\mathcal{L}$ after line 16. Then, there exists a path $\pi\in\Pi$ such that $\delta_{\pi}(\pi(S))<\delta_{\pi}(u)$, where $u$ is the unverified vertex along $\pi$.
@@
Since $S$ was put in $\mathcal{L}$, there must exist a vertex $s\in S$ such that the algorithm explored all bridges of $\Pi$ attached to $s$. This is guaranteed by lines 11 and 12. This implies that $s$ precedes the unverified vertex along its path. ∎
Next we prove that candidates that are not straddled by any bridges of $\Pi$ are $k$-shredders. This can be viewed as a local version of one direction of Lemma 4.8.
@@
Let $S$ be a set put in $\mathcal{L}$ at some point on line 13. If no bridge of $\Pi$ straddles $S$, then $S$ is a $k$-shredder.
@@
We prove $S$ is a $k$-shredder by showing $G\setminus S$ has at least three components. Firstly, we have $S=N(Q)$ for some component $Q$ of $G\setminus\Pi$. Fix a path $\pi\in\Pi$ such that the vertex $\pi(S)$ is not the far-most endpoint of $\pi$. Such a path must exist by 5.8. Let $z$ denote the far-most endpoint of $\pi$. It suffices to show that $x$ is not connected to $z$ in $G\setminus S$. Then, because $x\notin Q$ and $z\notin Q$, there are at least three components of $G\setminus S$: the component containing $x$, the component containing $z$, and $Q$.
@@
Consider the open subpath $P(u{\rightsquigarrow}v)$ (see Figure 4). We claim that every vertex in $P(u{\rightsquigarrow}v)$ is not in $\Pi$. To see this, suppose that there exists a vertex $w\in P(u{\rightsquigarrow}v)$ such that $w\in\pi_{3}$ for some path $\pi_{3}\in\Pi$. Then, $\delta_{\pi_{3}}(w)=\delta_{\pi_{3}}(\pi_{3}(S))$ because otherwise, it would contradict the definition of $u$ or $v$. This implies $w=\pi_{3}(S)$ because the vertex distance $\delta_{\pi_{3}}(\pi_{3}(S))$ from $x$ along $\pi_{3}$ is unique. However, this contradicts the definition of $P$ because $P$ is a path in $G\setminus S$: it cannot contain $\pi_{3}(S)$.
@@
Let $S$ be a candidate $k$-shredder detected by Algorithm 1 on input $(x,\nu,\Pi)$. If there exists a bridge of $\Pi$ that straddles $S$, then $S$ will be pruned from the list of candidates. Similarly, if there exists a bridge that straddles the unverified set $U$, then $U$ will be pruned.
@@
Let $S$ be the candidate straddled by a bridge $\Gamma$ of $\Pi$. For a visual example of this proof, see Figure 5. Let $A$ denote the attachment set of $\Gamma$. If any vertex in $A$ has been fully processed without terminating early, then $\Gamma$ will be detected on line 9. In this case, $S$ will be pruned by line 16.
Hence, we can assume that no vertex of $A$ has been fully explored, i.e. $U\preceq A$. We will show that $U$ straddles $S$, in which case $S$ will be pruned by line 16. First, by 5.8, there must exist a path $\pi_{1}\in\Pi$ that satisfies $\delta_{\pi_{1}}(\pi_{1}(S))<\delta_{\pi_{1}}(\pi_{1}(U))$. Because $\Gamma$ straddles $S$, there must exist a path $\pi_{2}\in\Pi$ and a vertex $v\in A\cap\pi_{2}$ such that $\delta_{\pi_{2}}(v)<\delta_{\pi_{2}}(\pi_{2}(S))$. Furthermore, because $U\preceq A$, this implies that $\delta_{\pi_{2}}(\pi_{2}(U))\leq\delta_{\pi_{2}}(v)<\delta_{\pi_{2}}(\pi_{2}(S))$. Now, we have $\delta_{\pi_{1}}(\pi_{1}(S))<\delta_{\pi_{1}}(\pi_{1}(U))$ and $\delta_{\pi_{2}}(\pi_{2}(U))<\delta_{\pi_{2}}(\pi_{2}(S))$, which implies $U$ straddles $S$. Line 16 implies that $S$ will be pruned from the list of candidates.
Figure 5: The bridge $Q$ has attachment set $A=S=\{s_{1},s_{2},s_{3}\}$. $S$ is straddled by $\Gamma$ with attachments $\{u,v\}$. If we explore all bridges of $\Pi$ attached to $u$, then $\Gamma$ will be naturally detected (top). Otherwise, we must have terminated early at some vertex $u_{1}$ before reaching $u$. 5.8 implies all bridges of $\Pi$ attached to some vertex $s_{2}$ were explored without terminating early. So $s_{2}$ must precede some unverified vertex $u_{2}$. This implies that $S$ is straddled by the unverified vertices (bottom).
Next, we prove the second statement. Suppose the unverified set $U$ is straddled by a bridge. First, we may assume that $|U|=k$ as otherwise $U=\{x\}$, which is caught by line 13. Suppose that a bridge $\Gamma$ of $\Pi$ straddles $U$ with attachment set $A$. If any vertex in $A$ has been fully explored by the algorithm, then $\Gamma$ will be detected by the BFS and $U$ will be pruned by line 16. Otherwise, we have $U\preceq A$, which contradicts our assumption that $\Gamma$ straddles $U$. ∎
Proof of Lemma 5.6.
Lemma 5.9 shows that the sets in $\mathcal{L}$ which are not straddled by a bridge of $\Pi$ are $k$-shredders. Lemma 5.10 shows that sets in $\mathcal{L}$ (along with the unverified set) which are straddled are removed from $\mathcal{L}$. This implies that at the end of Algorithm 1, every set in $\mathcal{L}$ is a $k$-shredder. ∎
Finally, we can combine Lemmas 5.5, 5.6 and 5.4 to prove Lemma 2.5.
Proof of Lemma 2.5.
Let $S$ be a $k$-shredder captured by $(x,\nu,\Pi)$. Lemma 5.5 implies that $S\in\mathcal{L}$ or $S=U$ by the end of the algorithm. Furthermore, every set in $\mathcal{L}$ by the end of the algorithm is a $k$-shredder by Lemma 5.6. Finally, Lemma 5.4 proves the time complexity of the algorithm. ∎